OpenStudy (anonymous):

Need help quickly please! Find the rectangular coordinates of the point with the polar coordinates. (3,2pi/3)

OpenStudy (sidsiddhartha):

To convert a point in polar coordinates (r, theta), use the conversion to rectangular coordinates (x, y) where x = rcos(theta) and y = rsin(theta).

OpenStudy (anonymous):

I don't know how to do that... could you show me?

OpenStudy (anonymous):

i have answer choices if that helps?

OpenStudy (sidsiddhartha):

x = r * cos(theta) = 3 * cos(2pi/3) y = r * sin(theta) = 3* sin(2pi/3)

OpenStudy (anonymous):

oh ok, now what?

OpenStudy (sidsiddhartha):

x,y are the rectangular coordinates just put the values of cos and sin

OpenStudy (anonymous):

that doesn't come to one of the options though

OpenStudy (anonymous):

(-3/2,3√3/2) (3√3/2,-3/2) (-3/2,3/2) (3/2,-3/2)

OpenStudy (sidsiddhartha):

cos2pi/3=-1/2 and sin 2pi/3=sqrt(3)/2 now can u do it?

OpenStudy (sidsiddhartha):

so what's ur answer?

OpenStudy (anonymous):

C?

OpenStudy (sidsiddhartha):

\[(-3/2,3\sqrt{3}/2)\]

OpenStudy (anonymous):

ohhhh ok thank you very much!

OpenStudy (sidsiddhartha):

np:)